# Evaluating Cooling-Center Coverage Using Persistent Homology of a Filtered Witness Complex

### v2

@@ 4.2. Topological Approach
We compute the PH of the filtered witness complex for each of the four locations of interest. Their persistence diagrams (PDs) are shown in Figure 4. Notably, the PD for Boston shows significantly more 1D homology class points than any other location. Figure 5 plots the distributions of 0D and 1D death values for the four locations of interest. The average death times (which are listed in Table 1) are fairly similar across cities in both dimension 0 and 1; the average death time in dimension 0 is between 0.5 and 2 while the average in dimension 1 is between 1 and 2.5. It is worth noting that Austin has many more outliers with high 0-dimensional death time than any other city, possibly indicating more regions with greater risk.
@@
For dimension 0, all features have a birth time of 0 so the death time is also the lifetime. For dimension 1, however, the birth time is nonzero so we also compute the lifetime of the persistence points. The birth time alone is not informative in the context of this application, however if the lifetime is small, it indicates that the hole fills in soon after it is formed so it may not be a significant hole in coverage. Even though in dimension 1 Austin has the second highest average death time, the average lifetime is the smallest across all four cities, likely indicating the regions indicated by the 1-dimensional death simplices are less vulnerable in Austin than in other cities. On the other hand, Boston has the largest death times and lifetimes of 1-dimensional features across all four cities, likely indicating there are more significant holes in coverage in Boston than in Austin.
@@
Figure 7. Death simplices for the 0-dimensional homology classes. Maps are not to scale and the scale of death values are different across the cities.
4.2.1. Death Simplices
@@
In dimension 1, the death simplices indicate the epicenter of the hole in coverage of cooling centers. In Figure 8, we see one or more death simplices with a high death time in the southern regions of Boston and Portland as well as in central Miami. These regions are the most vulnerable to heat related mortality. Boston also has numerous death simplices with greater-than-average death values. This indicates Boston has more significant holes in coverage than the other three cities.
Figure 8. Death simplices for the 1-dimensional homology classes. Largest death values are represented in yellow. Maps are not to scale.
Figure 9. Comparison of the two measures of vulnerability. The vertical axis plots the maximum death time of components containing any census block within a given census tract. The horizontal axis plots the corresponding HVI for the census tract. Table 2 shows the correlation and corresponding $p$-value for each city individually as well as all four combined.
Figure 10. In each map, census tracts are colored based on their HVI value relative to the average HVI within the city. Therefore pink and green tracts have an HVI above and below average, respectively. The blue lines and triangles are the death simplices for dimension 0 and 1 with death times in the 90% percentile of death times within that city.
4.3. Comparing Approaches
@@
Next we want to look at specific regions. To do so, we plot the death simplices with a death time in the 90-th percentile and the HVI data on the same plot in Figure 10. It is important to note the colorbar in Figure 10 does not represent the HVI. Instead, for a given census tract $T$, it plots $HVI(T)-\text{mean}\{HVI(t):t\in\mathcal{T}\}$ where $\mathcal{T}$ is the set of all census tracts within a given city. Therefore the colorbars should be interpreted as a measure of how vulnerable a tract is relative to the average. Census tracts colored pink (which corresponds to positive values) indicate census tracts with an HVI above the mean. Therefore, any intersection between a death simplex and a pink (above average HVI) census tract indicate regions of agreement between the two measures of vulnerability. We will focus on analyzing this plot to compare when the two methods align and analyze when they differ.
In Austin, we see significant overlap of the death simplices for dimension 0 with the regions of above-average HVI. In particular, on the northeastern portion of Austin, there are two census tracts with well above average HVI that also intersect with death simplices. This indicates agreement between the approaches meaning the regions with high HVI also lack cooling center coverage. Because both methodologies agree in these locations, we can interpret these locations a being more severely vulnerable than the locations identified by only one approach.
@@
We can also use this plot to identify regions of disagreement where one approach may identify the region as vulnerable while the other does not. These regions would indicate a moderate level of vulnerability For example, regions with above average HVI but no death simplices could be due to a significant vulnerable population but in an area that is well covered by cooling centers. Regions with a below average HVI but several death simplices could indicate areas that have lower temperatures and more tree cover but do not have many nearby cooling centers. An HVI well below average with death simplices might indicate a low-population area where cooling centers are therefore not as necessary. We see examples of all of these in Figure 10.
@@ 5.1. Limitations
Further, as discussed in , we use the geodesic distance between census blocks and cooling centers. While some work suggests this is an appropriate proxy for travel distance, it cannot account for variables that influence travel such as traffic and road closures. Ideally we would like to use travel time like the approach in [23], however google maps has a fee for each distance calculated, so using travel time or different modes of transportation was infeasible in this case.
Lastly, it is important to address the topographical features that may limit our analysis. For one, census blocks can be different sizes within a given city. This is a problem, for example, because a centroid in a large census block may appear to be isolated spatially in our analysis and, therefore, not connect to other centroids until later in the filtration. This census block could be labeled as not being well covered, when in fact it is. Ideally, a census block’s centroid would be a good geographical approximation for any neighborhood within the block. A future improvement could include sampling multiple points within a census block rather than only using the centroid. In addition, our analysis does not account for bodies of water. For example, some unpopulated islands off the coast of Boston are included within census block boundaries without regards to the ocean.
Table 2. Correlation between death time and HVI, as shown in Figure 9. The last row labeled “All” indicates the correlation if we combined the points for all four cities. Note that Boston is the only city with a $p$-value <0.05<0.05 with a correlation of 0.132, indicating a slightly positive correlation.
City Correlation Coeff. p-value
Austin 0.096 0.446
Boston 0.132 0.018
Miami -0.140 0.171
Portland -0.075 0.372
All 0.048 0.234
5.2. Future Work
In the future, we would like to explore other spatially aware approaches for detecting heat vulnerability from the literature as additional points of comparison. We’d also like to conduct a larger scale investigation about additional cities. We chose four different cities to test our method, however there are many other cities included in the US Urban Heat Island Mapping Campaign.
@@ Appendix A Data Collection
In this section we describe the data sources along with any pre-processing steps applied to them. A.1 describes the landmark (census block centroids) data and A.2 describes the witness (cooling center) data. Figure 11 plots the full dataset of landmarks and witnesses.
Figure 11. Final dataset of census tract centroids (landmarks) and cooling centers (witnesses) for each of the four locations, colored in light green and dark green, respectively.
A.1. Landmark Data Collection and Processing
@@
A.2. Witness Data Collection and Processing
@@
Table 3. Furthest Southwest and Northeast coordinates used to search OpenStreetMap for cooling center tags.
@@
A.3. Heat Vulnerability Index Data Collection and Processing
@@
Data on the typical afternoon temperature, the percentage of the area covered in tree canopy, number of individuals who are under 5, and number of individuals who are older than 65 was retrieved from the National Integrated Heat Health Information System (NIHHIS). The dataset was created as part of the National Oceanic and Atmospheric Administration’s (NOAA’s) “U.S. Urban Heat Island Mapping Campaign.” Data on the area not covered in tree canopy was found by subtracting the data on the percent of area covered in tree canopy in a given census tract from 100. Figures 1215 show the the individual HVI variables for each city.
Our motivation for incorporating these particular four variables into our measure of vulnerability is detailed as follows. Firstly, one can imagine that a warmer census block with few trees may not fare well during an extreme heat event. Secondly, being over the age of 65 puts an individual more at risk of heat related mortality due to social isolation and a lack of mobility to get assistance when faced with emergencies [19]. Lastly, because children thermoregulate less effectively than adults, they are more likely to respond poorly during extreme heat events [19]. The census block is interpreted to be at greater risk when its HVI is higher. See Figure 3 for the maps of the HVI across the four regions of interest. More details on the demographic data used to create the HVI can be found in Appendix A.3.
@@ Appendix B Additional Data Analysis
We tested the independence of the four variables from the HVI using variance inflation factor (VIF). The results are presented in Table 4. These results are discussed in Section 3.2. We further explore the relationships between the HVI variables by considering their pairwise correlations. Figures 1619 show scatterplots of all pairwise relationships and the corresponding As expected, from the VIF analysis, Boston and Portland have a strong pairwise correlation between the typical afternoon temperature and the percentage of the area not covered with tree canopy. In Portland, however, neither variable is strongly correlated with either demographic variables.
Austin
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 12. Heatmaps of the four HVI variables for Austin. Top left: area not covered by tree canopy; top right: typical afternoon temperature; bottom left: number of individuals of age less than 5 years old; bottom right: number of individuals of age greater than 65.
Boston
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 13. Heatmaps of the four HVI variables for Boston. Top left: area not covered by tree canopy; top right: typical afternoon temperature; bottom left: number of individuals of age less than 5 years old; bottom right: number of individuals of age greater than 65.
Miami
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 14. Heatmaps of the four HVI variables for Miami. Top left: area not covered by tree canopy; top right: typical afternoon temperature; bottom left: number of individuals of age less than 5 years old; bottom right: number of individuals of age greater than 65.
Portland
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 15. Heatmaps of the four HVI variables for Portland. Top left: area not covered by tree canopy; top right: typical afternoon temperature; bottom left: number of individuals of age less than 5 years old; bottom right: number of individuals of age greater than 65.
Table 4. Variance inflation factor (VIF) results to check multicollinearity in the four cities of interest. Note: a VIF of 1 indicates no multicollinearity. VIFs between 1 and 3 are considered to have negligible to minimal multicollinearity [22].
@@
Table 5. The mean, median, and standard deviation (SD) of the HVI values across all census tracts within a given city.
City Mean Median SD
Austin 7.819 7.427 1.746
Boston 7.326 7.26 1.499
Miami 8.715 8.182 2.051
Portland 7.046 7.076 1.805
Figure 16. (a) Histograms and pairwise scatter plots of HVI variables across census tracts in Austin; (b) Pairwise Pearson correlations between the specified variables. A star indicates the correlation is not statistically significant as it does not have a $p$-value of $p<0.05$.
Figure 17. (a) Histograms and pairwise scatter plots of HVI variables across census tracts in Boston; (b) Pairwise Pearson correlations between the specified variables. A star indicates the correlation is not statistically significant as it does not have a $p$-value of $p<0.05$.
Figure 18. (a) Histograms and pairwise scatter plots of HVI variables across census tracts in Miami; (b) Pairwise Pearson correlations between the specified variables. A star indicates the correlation is not statistically significant as it does not have a $p$-value value of $p<0.05$.
Figure 19. (a) Histograms and pairwise scatter plots of HVI variables across census tracts in Portland; (b) Pairwise Pearson correlations between the specified variables. A star indicates the correlation is not statistically significant as it does not have a $p$-value of $p<0.05$.

### v3

@@ 4.2 Topological Approach
We compute the PH of the filtered witness complex for each of the four locations of interest. Their persistence diagrams (PDs) are shown in Figure 4. Notably, the PD for Boston shows more 1D homology class points than any other location. Figure 5 plots the distributions of 0D and 1D death values for the four locations of interest. The average death times (which are listed in Table 1) are fairly similar across cities in both dimension 0 and 1; the average death time in dimension 0 is between 0.5 and 2 while the average in dimension 1 is between 1 and 2.5. It is worth noting that Austin has many more outliers with high 0-dimensional death time than any other city, possibly indicating more regions with greater risk.
@@
For dimension 0, all features have a birth time of 0 so the death time is also the lifetime. For dimension 1, however, the birth time is nonzero so we also compute the lifetime of the persistence points. The birth time alone is not informative in the context of this application, however if the lifetime is small, it indicates that the hole fills in soon after it is formed so it may not be a notable hole in coverage. Even though in dimension 1 Austin has the second highest average death time, the average lifetime is the smallest across all four cities, likely indicating the regions indicated by the 1-dimensional death simplices are less vulnerable in Austin than in other cities. On the other hand, Boston has the largest death times and lifetimes of 1-dimensional features across all four cities, likely indicating there are more notable holes in coverage in Boston than in Austin.
@@
Figure 7: Death simplices for the 0-dimensional homology classes. The black dots represent the locations of cooling centers, including those outside the city but in neighboring areas. Maps are not to scale and the scale of death values are different across the cities.
4.2.1 Death Simplices
@@
In dimension 1, the death simplices indicate the epicenter of the hole in coverage of cooling centers. In Figure 8, we see one or more death simplices with a high death time in the southern regions of Boston and Portland as well as in central Miami. These regions are the most vulnerable to heat related mortality. Boston also has numerous death simplices with greater-than-average death values. This indicates Boston has more notable holes in coverage than the other three cities.
Figure 8: Death simplices for the 1-dimensional homology classes. The black dots represent the locations of cooling centers, including those outside the city but in neighboring areas. Largest death values are represented in yellow. Maps are not to scale.
Figure 9: Comparison of the two measures of vulnerability. The vertical axis plots the maximum death time of components containing any census block within a given census tract. The horizontal axis plots the corresponding HVI for the census tract. Table 2 shows the correlation and corresponding $p$-value for each city individually as well as all four combined.
Figure 10: In each map, census tracts are colored based on their HVI value relative to the average HVI within the city. Therefore pink and green tracts have an HVI above and below average, respectively. The blue lines and triangles are the death simplices for dimension 0 and 1 with death times in the 90% percentile of death times within that city.
4.3 Comparing Approaches
@@
Next we want to look at specific regions. To do so, we plot the death simplices with a death time in the 90-th percentile and the HVI data on the same plot in Figure 10. It is important to note that the colorbar in Figure 10 does not represent the HVI. Instead, for a given census tract $T$, it plots $HVI(T)-\text{mean}\{HVI(t):t\in\mathcal{T}\}$ where $\mathcal{T}$ is the set of all census tracts within a given city. Therefore the colorbars should be interpreted as a measure of how vulnerable a tract is relative to the average. Census tracts colored pink (which corresponds to positive values) indicate census tracts with an HVI above the mean. Therefore, any intersection between a death simplex and a pink (above average HVI) census tract indicate regions of agreement between the two measures of vulnerability. We will focus on analyzing this plot to compare when the two methods align and analyze when they differ.
In Austin, we see overlap of the death simplices for dimension 0 with the regions of above-average HVI. In particular, on the northeastern portion of Austin, there are two census tracts with well above average HVI that also intersect with death simplices. This indicates agreement between the approaches meaning the regions with high HVI also lack cooling center coverage. Because both methodologies agree in these locations, we can interpret these locations a being more severely vulnerable than the locations identified by only one approach.
@@
We can also use this plot to identify regions of disagreement where one approach may identify the region as vulnerable while the other does not. These regions would indicate a moderate level of vulnerability. For example, regions with above average HVI but no death simplices could be due to a high vulnerable population but in an area that is well covered by cooling centers. Regions with a below average HVI but several death simplices could indicate areas that have lower temperatures and more tree cover but do not have many nearby cooling centers. An HVI well below average with death simplices might indicate a low-population area where cooling centers are therefore not as necessary. We see examples of all of these in Figure 10.
@@ 5.1 Limitations
Further, as discussed in Section 3.1.2, we use the geodesic distance between census blocks and cooling centers. While some work suggests this is an appropriate proxy for travel distance, it cannot account for variables that influence travel such as traffic and road closures. Ideally we would like to use travel time like the approach in [24], however google maps has a fee for each distance calculated, so using travel time or different modes of transportation was infeasible in this case.
It is also important to address the topographical features that may limit our analysis. For one, census blocks can be different sizes within a given city. This is a problem, for example, because a centroid in a large census block may appear to be isolated spatially in our analysis and, therefore, not connect to other centroids until later in the filtration. This census block could be labeled as not being well covered, when in fact it is. Ideally, a census block’s centroid would be a good geographical approximation for any neighborhood within the block. A future improvement could include sampling multiple points within a census block rather than only using the centroid. In addition, our analysis does not account for bodies of water. For example, some unpopulated islands off the coast of Boston are included within census block boundaries without regards to the ocean.
Lastly, we note again that death simplices are not stable. Thus, small changes in the data could change the location of the death simplex, leading to different conclusions.
Table 2: Correlation between death time and HVI, as shown in Figure 9. The $p$-value is computed from a two-sided $t$-test. The last column labeled “All” indicates the correlation if we combined the points for all four cities. Note that Boston is the only city with a $p$-value <0.05<0.05 with a correlation of 0.132, indicating a slightly positive correlation.
Austin Boston Miami Portland All
Correlation Coefficient 0.096 0.132 -0.140 -0.075 0.048
p-value 0.446 0.018 0.171 0.372 0.234
5.2 Future Work
In the future, we would like to improve our method by changing how we capture the vulnerable regions from the topological approach. As discussed, death simplices are not stable which can lead to different results with small changes in the input data. Using minimal cycle representatives [31] to identify the holes in coverage could avoid this issue and allow for better identification of areas with poor access to resources. However, this raises numerous other decisions and concerns as minimal cycle representatives are not uniquely defined and can be challenging to compute.
We would also like to explore other spatially aware approaches for detecting heat vulnerability from the literature as additional points of comparison. We’d also like to conduct a larger scale investigation about additional cities. We chose four different cities to test our method, however there are many other cities included in the US Urban Heat Island Mapping Campaign.
@@ Appendix A Data Collection
In this section we describe the data sources along with any pre-processing steps applied to them. Section A.1 describes the landmark (census block centroids) data, Section A.2 describes the witness (cooling center) data, and Section A.3 describes the HVI data. Figure 11 plots the full dataset of landmarks and witnesses.
Figure 11: Final dataset of census tract centroids (landmarks) and cooling centers (witnesses) for each of the four locations, indicated as black dogs and blue diamonds, respectively.
A.1 Landmark Data Collection and Processing
@@
A.2 Witness Data Collection and Processing
@@
Table 3: Furthest Southwest and Northeast coordinates used to search OpenStreetMap for cooling center tags.
@@
A.3 Heat Vulnerability Index Data Collection and Processing
@@
Data on the typical afternoon temperature, the percentage of the area covered in tree canopy, number of individuals who are under 5, and number of individuals who are older than 65 was retrieved from the National Integrated Heat Health Information System (NIHHIS). The dataset was created as part of the National Oceanic and Atmospheric Administration’s (NOAA’s) “U.S. Urban Heat Island Mapping Campaign.” Data on the area not covered in tree canopy was found by subtracting the data on the percent of area covered in tree canopy in a given census tract from 100. Figure 12 shows the the individual HVI variables for each city.
Our motivation for incorporating these particular four variables into our measure of vulnerability is detailed as follows. Firstly, one can imagine that a warmer census block with few trees may not fare well during an extreme heat event. Secondly, being over the age of 65 puts an individual more at risk of heat related mortality due to social isolation and a lack of mobility to get assistance when faced with emergencies [19]. Lastly, because children thermoregulate less effectively than adults, they are more likely to respond poorly during extreme heat events [19]. The census block is interpreted to be at greater risk when its HVI is higher.
@@ Appendix B Additional Data Analysis
We tested the independence of the four variables from the HVI using variance inflation factor (VIF). The results are presented in Table 4. These results are discussed in Section 3.2.1. We further explore the relationships between the HVI variables by considering their pairwise correlations. Figures  13 and 14 and show scatterplots of all pairwise relationships and the corresponding correlation coefficients. As expected, from the VIF analysis, Boston and Portland have a strong pairwise correlation between the typical afternoon temperature and the percentage of the area not covered with tree canopy. In Portland, however, neither variable is strongly correlated with either demographic variables.
Table 4: Variance inflation factor (VIF) results to check multicollinearity in the four cities of interest. Note: a VIF of 1 indicates no multicollinearity. VIFs between 1 and 3 are considered to have negligible to minimal multicollinearity [23].
@@
Austin
Refer to caption Refer to caption Refer to caption Refer to caption
Boston
Refer to caption Refer to caption Refer to caption Refer to caption
Miami
Refer to caption Refer to caption Refer to caption Refer to caption
Portland
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 12: Heatmaps of the four HVI variables for Austin, Boston, Miami, and Portland. From left to right: area not covered by tree canopy, typical afternoon temperature, number of individuals of age less than 5 years old, and number of individuals of age greater than 65.
Figure 13: Histograms and pairwise scatter plots of HVI variables across census tracts within each city.
Figure 14: Matrices of pairwise Pearson correlations between the specified variables within each city. A star indicates the correlation is not statistically significant according to the two-sided $t$-test as it does not have a $p$-value of $p<0.05$.